# From Roots to Rewards: Dynamic Tree Reasoning with Reinforcement Learning

### v3

@@ Abstract
Index Terms:
Reinforcement learning, Tree-of-Thought, Question answering, Large language models, Probabilistic reasoning.
@@ II Reinforcement Learning Framework for Dynamic Tree-of-Thought Reasoning
To overcome the limitations of static and brute-force tree generation in the original ProbTree framework [1], we introduce a reinforcement learning (RL) based agent that dynamically generates reasoning trees and selects actions at each step to answer knowledge-intensive questions efficiently and accurately.
@@ II-B Dynamic Tree Construction via RL Agent
Figure 1: RL agent-guided reasoning. At each node, the agent selects one action among CB, OB, or Child based on predicted probabilities. In this example, the Child action has the highest probability (50%) and is selected. CB and OB (20%, 30%) are not executed. CB/OB would return answers directly, whereas Child expands the tree, recurses over new child nodes, and aggregates intermediate answers.
Legend: Blue = Data/Input, Orange = Model/Decision, Green = Reasoning Process
@@ II-C Deep Q-Learning Architecture
Reward Function:
Given a chosen answer $a_{\text{chosen}}$, gold answer $a_{\text{gold}}$, and LLM call cost $c$, the reward is computed as:
@@ II-E Training Regimes
Each variant was trained under multiple reward configurations:
@@ III Experimental Setup
In this section, we evaluate the performance of our proposed models across three datasets: HotpotQA, Musique, and 2WikiMultihopQA. We consider different training and testing configurations to analyze generalization, efficiency, and robustness.
@@ III-A Datasets
III-B Evaluation Metrics
We assess performance using:
  • Accuracy: Percentage of correctly predicted answers that match gold answers (verified by asking the LLM if both answers match).
  • LLM Call Count: Total LLM invocations during inference.
III-C Training and Testing Configurations
@@ IV Evaluation and Results
IV-A LLM Setup
@@
IV-B In-domain Results
IV-B1 In-domain Accuracy Comparison Across Datasets
@@
TABLE II: In-domain accuracy (%) across reward configurations and datasets. Columns labeled ”Orig.” represent the results from models in their ”Original Configurations” (e.g., Original ProbTree, Greedy Solver, Random Agent), which do not use the reward setups (High, Balanced, Efficient). Remaining columns represent RL-based models under the reward configurations. The best accuracy for each dataset (across Original Config and RL-based Configs) is bolded. - indicates configurations that are not applicable.
@@
IV-B2 Accuracy-Cost Tradeoff Analysis
@@
Figure 3: Accuracy vs. Number of LLM calls tradeoff across datasets, models, and configurations. Evaluation was conducted on a set of 100 test examples per dataset.
@@
IV-C Results: Cross-Domain Generalization
@@ 36.60 Resampling Balanced
IV-D Results: Multi-Domain Generalization
@@
TABLE IV: Multi-Domain Generalization Results: Accuracy (%) across Datasets and Approaches. Bold values indicate the best performance within each dataset.
Approach / Configuration HotpotQA 2Wiki Musique
Original ProbTree 71.40 74.60 44.80
Greedy Solver 70.80 77.20 45.20
RL (Question Only)
High Accuracy 58.60 55.20 25.20
@@
RL (Question + CB + OB)
High Accuracy 60.00 55.20 24.80
@@
RL (Question + CB + OB + Reformulation)
High Accuracy 61.80 58.80 25.80
@@
RL (Transformer, Question Only)
High Accuracy 69.20 71.60 21.60
@@
RL (Resampling)
High Accuracy 67.80 74.00 41.60
@@
IV-E Method Usage Analysis
@@ V-A Structured Prompting and Reasoning Topologies
Recent advancements in structured prompting have introduced various reasoning topologies to enhance the capabilities of large language models (LLMs). Besta et al. [12] provide a comprehensive taxonomy of these strategies, categorizing them along three dimensions: Topology Scheme, Reasoning Schedule, and AI Pipeline involvement.
Topology Scheme: This dimension classifies the structural format of reasoning, such as chains [13], trees [14], or graphs. Our approach aligns with the Tree-of-Thoughts [14] paradigm, where reasoning is structured hierarchically. However, unlike static trees, our method employs a dynamic tree construction guided by a reinforcement learning (RL) agent, allowing for adaptive reasoning paths based on the problem context.
@@ V-B Adaptive Reasoning Frameworks
V-C Self-Guided Multimodal Reasoning
In the realm of multimodal reasoning, self-guided strategies have been explored to enhance LLM performance. Hu et al. [16] propose the Socratic Questioning framework, where the model engages in self-questioning to guide its reasoning process, particularly in multimodal contexts. This approach heuristically guides multimodal LLMs to focus on relevant visual clues, reducing hallucinations and enhancing the model’s ability to describe fine-grained image details.
While Socratic Questioning emphasizes recursive self-inquiry to improve reasoning depth, our method focuses on strategic action selection through reinforcement learning to optimize both reasoning quality and resource utilization. Both approaches aim to enhance the reasoning capabilities of LLMs, but they operate in different modalities and utilize distinct mechanisms for guiding the reasoning process.
V-D Positioning Our Approach
@@ VI Conclusion
We presented a reinforcement learning-based framework for dynamic Tree-of-Thought (ToT) reasoning that enables large language models (LLMs) to perform adaptive, multi-step decision-making during inference. By formulating reasoning as a Markov Decision Process, the framework allows the agent to strategically select actions such as decomposition, retrieval, or direct answering, depending on the task’s complexity and context. This dynamic construction of reasoning trees balances efficiency and depth, offering significant improvements in both accuracy and computational resource utilization.
@@
Appendix A Complete Cross-Domain Generalization Results
Table V presents the full results across all configurations and feature sets for cross-domain generalization experiments. Since the Greedy configuration achieves the best performance across all cases, we highlight it in bold. Additionally, for comparison, we also bold the second-best result among the RL-based configurations (excluding Greedy) for each train-test pair to indicate the next best-performing model in terms of accuracy.
TABLE V: Detailed cross-domain generalization results: Accuracy (%) across train-test pairs. Rows include all feature sets and configurations.
Model + Config Train → Test Datasets
Musique Musique 2Wiki 2Wiki HotpotQA HotpotQA
HotpotQA 2Wiki Musique HotpotQA 2Wiki Musique
Greedy 69.20 73.00 46.00 70.60 74.40 44.20
Q-Only (High Accuracy) 59.20 59.20 24.80 59.40 49.00 23.20
Q-Only (Balanced) 60.60 60.60 26.00 59.40 49.00 24.20
Q-Only (Efficiency) 61.00 61.00 24.80 56.60 46.80 27.00
Q+CB+OB (High Accuracy) 60.80 52.80 25.60 59.60 56.20 26.60
Q+CB+OB (Balanced) 61.20 55.00 25.40 60.40 57.60 36.00
Q+CB+OB (Efficiency) 61.80 56.20 27.20 57.40 57.60 36.60
Transformers DQN (High Accuracy) 59.20 52.20 23.20 55.80 71.60 21.60
Transformers DQN (Balanced) 57.60 52.80 26.80 62.20 57.00 27.00
Transformers DQN (Efficiency) 63.00 57.00 23.20 55.80 54.00 27.20
Reformulation (High Accuracy) 61.20 54.40 25.00 58.40 57.60 24.40
Reformulation (Balanced) 58.60 55.20 26.20 59.80 53.40 24.60
Reformulation (Efficiency) 59.80 57.00 27.20 60.40 53.80 27.00
Resampling (High Accuracy) 57.20 62.40 32.20 63.40 63.40 34.40
Resampling (Balanced) 60.80 60.40 36.60 59.40 62.00 36.60
Resampling (Efficiency) 58.60 65.20 34.00 58.80 55.80 36.40
Appendix B Method Usage by Model Variant
While the main paper discusses method usage across reward configurations, we complement that view by analyzing method preferences by model architecture. Figure 5 shows the distribution of selected reasoning strategies for each model variant across all datasets.
Figure 5: Method usage distribution per model variant across datasets. Each bar reflects the number of times a reasoning method (CB, OB, Child, etc.) was selected.
The breakdown reveals several key observations:
  • Question-Only agents rely heavily on CB and OB, with occasional Child selections, defaulting to simpler strategies.
  • Question + CB + OB models show a similar pattern but with slightly more Child usage from improved state representations derived from answer log-probabilities.
  • Reformulation-based agents display a more balanced distribution across original and reformulated variants, with CB_REFORMULATE and OB_REFORMULATE used significantly more, indicating that the additional action space allows for more nuanced strategies.
  • Transformer-based models strongly prefer Child, especially in HotpotQA, aligning with their superior accuracy in that domain.
  • Resampling models are skewed toward the ResampleTreeDecompositionThenSolve action, confirming their tendency to revise poor decompositions to enhance performance.
This model-wise analysis complements the configuration-wise view and highlights how architectural choices shape the agent’s use of its action space.
Appendix C State Representations for Models
Table VI summarizes the features used to construct the state representations for each of our models. These state features include statistical attributes, semantic embeddings, structural information, and confidence scores. Each model includes a subset of the available features based on its specific design and configuration (e.g., Question Only, CB+OB, Resampling, etc.). Feature Descriptions. Below we summarize the intuition behind each feature group:
  • Has Children (0/1): Binary flag indicating whether a node can expand further, distinguishing leaves from intermediate states.
  • Question Length and Type: Captures surface attributes of the input (short vs. multi-hop), which affect decomposition depth.
  • Number of Children: The branching factor of a node, influencing search space size and pruning strategies.
  • Success Rates for Each Method: Empirical performance of decomposition methods (e.g., CB, OB) measured on the development set, used to guide policy decisions.
  • Question Embedding: Semantic vector of the question, supporting generalization across similar inputs.
  • Tree Depth and Position: Encodes node placement in the reasoning tree, indicating reasoning progress.
  • Log Probabilities: Likelihoods from the LM, used as uncertainty signals for exploration vs. exploitation.
  • Answer Embeddings: Semantic vectors of candidate answers, enabling consistency checks across branches.
  • Answer Lengths: Heuristic indicator, as very short or long outputs often correlate with lower quality.
  • Answer Confidence Scores: Certainty scores for candidate answers, key for pruning and selection.
  • Child-Specific Features: Per-child versions of the above signals (e.g., length, confidence, overlap), enabling fine-grained resampling and prioritization.
TABLE VI: State Feature Usage Across Models
State Features Question Only Transformers CB+OB Reformulation Resampling
Has Children (0/1)
Question Length and Type
Number of Children
Success Rates for each method
Question Embedding
Tree Depth and Position
LogProbabilities
Answer Embeddings
Answer Lengths
Answers Confidence Scores
Child-Specific Features

### v4

@@ II Reinforcement Learning Framework for Dynamic Tree-of-Thought Reasoning
To overcome the limitations of static and brute-force tree generation in the original ProbTree framework [1], we introduce a RL based agent that dynamically generates reasoning trees and selects actions at each step to answer knowledge-intensive questions efficiently and accurately.
@@ II-B Dynamic Tree Construction via RL Agent
Figure 1: RL agent-guided reasoning. The figure illustrates how the agent selects among CB, OB, or Child actions. Here, Child (50%) is chosen, expanding the reasoning tree, while CB (20%) and OB (30%) would instead return answers directly.
Legend: Blue = Data/Input, Orange = Model/Decision, Green = Reasoning Process.
@@ II-C Deep Q-Learning Architecture
Reward Function: Given a chosen answer $a_{\text{chosen}}$, gold answer $a_{\text{gold}}$, and LLM call cost $c$, the reward is computed as:
@@ II-E Training Regimes
Each variant was trained under different reward configurations to capture the trade-off between accuracy and efficiency. We experimented with multiple values of $\alpha$ and $\beta$ on the development sets and found the following settings to be most effective for their intended objectives:
@@ III Experimental Setup
In this section, we describe our experimental setup, including the retrieval mechanism, evaluation metrics, and training/testing configurations. Our goal is to analyze model performance in terms of accuracy, efficiency, and generalization across multiple multi-hop QA benchmarks.
@@ III-A Datasets
III-B Retriever Details
For all experiments, we adopt the retrieval component from the original ProbTree implementation [1]. Following IRCoT [12], we use BM25 [13] implemented via Elasticsearch as our retriever. For each node $q_{i}\in T$, we retrieve the top-K paragraphs $R(q_{i})$, with $K\in\{3,5,7\}$ chosen based on development set performance. No additional reranking strategies are applied.
III-C Evaluation Metrics
We assess performance using both semantic correctness and efficiency measures:
  • Accuracy: Percentage of predictions judged as correct by the LLM-as-Judge [14], where the judge model provides binary (0/1) assessments, capturing diverse valid outputs beyond traditional metrics (e.g., EM, F1).
  • LLM Call Count: Total number of LLM invocations during inference.
III-D Training and Testing Configurations
@@ IV Evaluation and Results
IV-A Baselines and Comparison Scope
We primarily compare our proposed models against the original ProbTree framework, which was already evaluated against a range of adaptive reasoning baselines, including IRCoT [12] and ReAct [15], among others, using the same datasets, retrievers, and evaluation settings [1]. Therefore, demonstrating improvements over ProbTree on our datasets implies improvements over these baselines as well. For clarity and reproducibility, we refer readers to the original ProbTree paper for detailed comparisons with these adaptive methods.
IV-B LLM Setup
@@
For reproducibility, we adopt the same prompt templates as ProbTree [1]. The exact prompts and additional qualitative examples are available in our project repository.
IV-C In-domain Results
IV-C1 In-domain Accuracy Comparison Across Datasets
@@
TABLE II: In-domain accuracy (%) across reward configurations and datasets. ”Orig.” denotes models in their original setup (e.g., ProbTree, Greedy, Random) without reward configurations. Other columns report RL-based models under High, Balanced, and Efficient rewards.
@@
IV-C2 Accuracy-Cost Tradeoff Analysis
@@
Figure 3: Accuracy vs. Number of LLM calls tradeoff across datasets, models, and configurations. This evaluation was conducted on a subset of 100 test examples per dataset, specifically for analyzing efficiency.
@@
IV-D Results: Cross-Domain Generalization
@@ 36.60 Resampling Balanced
IV-E Results: Multi-Domain Generalization
@@
TABLE IV: Multi-Domain Generalization Results: Accuracy (%) across Datasets and Approaches.
Approach Configuration HotpotQA 2Wiki Musique
Original ProbTree - 71.40 74.60 44.80
Greedy Solver - 70.80 77.20 45.20
RL (Question Only) High Accuracy 58.60 55.20 25.20
@@
RL (Question + CB + OB) High Accuracy 60.00 55.20 24.80
@@
RL (Question + CB + OB + Reformulation) High Accuracy 61.80 58.80 25.80
@@
RL (Transformer, Question Only) High Accuracy 69.20 71.60 21.60
@@
RL (Resampling) High Accuracy 67.80 74.00 41.60
@@
IV-F Method Usage Analysis
@@ V-A Structured Prompting and Reasoning Topologies
Recent advancements in structured prompting have introduced various reasoning topologies to enhance the capabilities of LLMs. Besta et al. [16] provide a comprehensive taxonomy of these strategies, categorizing them along three dimensions: Topology Scheme, Reasoning Schedule, and AI Pipeline involvement.
Topology Scheme: This dimension classifies the structural format of reasoning, such as chains [17], trees [18], or graphs. Our approach aligns with the Tree-of-Thoughts [18] paradigm, where reasoning is structured hierarchically. However, unlike static trees, our method employs a dynamic tree construction guided by a RL agent, allowing for adaptive reasoning paths based on the problem context.
@@ V-B Adaptive Reasoning Frameworks
V-C Positioning Our Approach
@@ VI Conclusion
We presented a reinforcement learning-based framework for dynamic Tree-of-Thought (ToT) reasoning that enables LLMs to perform adaptive, multi-step decision-making during inference. By formulating reasoning as a Markov Decision Process, the framework allows the agent to strategically select actions such as decomposition, retrieval, or direct answering, depending on the task’s complexity and context. This dynamic construction of reasoning trees balances efficiency and depth, offering significant improvements in both accuracy and computational resource utilization.